# Accelerating Data Chunking in Deduplication Systems using Vector Instructions

### v1

@@ 2.3. Vector Instructions
Vector instructions (vector_instruction_sets, ) are special Single-Instruction Multiple-Data (SIMD) instructions supported by most modern processors. They rely on special vector registers for their operations. These registers come in multiple sizes; depending on the width of the vector register they use, vector instructions can be classified into different families (vector_instruction_sets, ). The most common vector register sizes are 128 bits, 256 bits, and 512 bits, i.e., 16, 32, and 64 bytes wide.
@@
Vector instructions support various arithmetic operations (intelIntrinsicsGuide, ), including pairwise addition, subtraction, multiplication, and maximum/minimum on packed values. Additionally, they support logical operations such as bitwise AND (&) and bitwise OR (|). They have been previously used to accelerate matrix multiplication (avx_matrixmul, ), sorting (avx_quicksort, ), multimedia applications (avx_multimedia, ), fluid simulations (avx_fluidmech, ), hash tables (hashtables_vector, ), and relational databases (fused_table_scan, ). The supported vector instruction types and their relative performance vary across CPU manufacturers.
@@ 2.3.1. Intel and AMD
Vector instructions on x86 platforms can be classified into three families: SSE-128, AVX-256, and AVX-512 (vector_instruction_sets, ). SSE-128 instructions use 128-bit registers and have been supported by Intel and AMD processors since 1999 (intelSIMDInstructions, ) and 2003 (AMD_SSE128, ), respectively. AVX-256 instructions were introduced by Intel and AMD in 2011 (intelSIMDInstructions, ), and use 256-bit registers. Finally, only a handful of the newest Intel (intelXeonPhi, ) and AMD (amdzen4_wikichip, ) processors, which have 512-bit wide vector registers, support AVX-512 instructions.
@@ 2.3.2. ARM
ARM processors have supported NEON-128 instructions, an equivalent to SSE-128, since 2011 (armNeonOverview, ). Modern ARM processors also support 256-bit and higher vector widths with the SVE / SVE2 instruction sets, available since 2021 (arm_sve, ). These two instruction sets differ in the kinds of instructions supported. For instance, NEON-128 does not support native VMASK operations, which are used to create integer masks by extracting one out of every k bits in a vector register, while SVE / SVE2 does. This can lead to performance differences in applications that need the VMASK operation (hashtables_vector, ).
@@ 2.3.3. IBM Power
IBM’s Power (ibm_power8, ) architecture supports AltiVec / VSX-128 vector instructions (altivec_instructions, ), an equivalent to SSE-128, since the 1990s. This instruction set supports equivalents for most SSE instructions but lacks VMASK support.
@@ 3.1. Performance bottlenecks in data deduplication
Figure 3. Time taken for Data Chunking vs Fingerprinting while deduplicating randomized data with FastCDC (fastcdc, ) and an 8 KB average chunk size.
Although both data chunking and fingerprinting have traditionally been considered the main bottlenecks in deduplication (dedup_intro, ), this has changed with the advent of new hashing algorithms/acceleration methods for fingerprinting. Figure 3 is a stacked bar plot showing the time taken by the data chunking and fingerprinting phases during deduplication. For fingerprinting, we use five different collision-resistant hashing algorithms (md5, ; sha1, ; sha256_512, ; appleby2008murmurhash, ; hashing_survey_xxhash, ). For data chunking, we use FastCDC (fastcdc, ), the fastest unaccelerated chunking algorithm (§6.2), with an 8 KB average chunk size. We use 30 GB of randomized data and an Intel Emerald Rapids machine described in §6 for this experiment.
Figure 3 shows that fingerprinting and data chunking take nearly equal time with traditional collision-resistant hashing algorithms, such as SHA1 (sha1, ) and SHA-256 (sha256_512, ). Fingerprinting takes longer than data chunking with other algorithms, such as MD5 (md5, ) and SHA-512 (sha256_512, ). This indicates that both phases used to be the performance bottlenecks in deduplication, aligning with previous literature (dedup_intro, ).
However, recent research (tunable_encrypted_dedup, ; scalable_incremental_checkpt, ) has introduced faster hashing algorithms such as MurmurHash3 (appleby2008murmurhash, ) and xxHash3 (xxhashWebsite, ). Figure 3 shows that fingerprinting takes significantly lower time than chunking with these new hashing algorithms, as they are $10\times-15\times$ faster than their counterparts on CPUs. Using GPUs can further accelerate fingerprinting speeds by up to $53\times$ (gpu_dedup, ; storegpu, ). Note that the choice of fingerprinting algorithm does not affect deduplication space savings, as MurmurHash3 and xxHash3 have similar collision resistance properties to other algorithms (appleby2016smhasher, ).
Thus, as a result of its computationally intensive nature and position on the critical path, data chunking is the main performance bottleneck in the deduplication pipeline and needs further optimization.
@@ 3.2. Accelerating hash-based algorithms with vector instructions
Figure 4. SS-CDC (sscdc, ): Accelerating the rolling hash phase
To address the data chunking bottleneck, SS-CDC (sscdc, ) proposed using AVX-512 instructions to accelerate hash-based CDC algorithms. They decouple the rolling hash and boundary detection phases, running the rolling hash on the entire source data to identify boundary candidates in the first phase, and determining boundaries sequentially in the second. This allows both stages to be independently accelerated with AVX instructions.
Figure 4 shows how SS-CDC (sscdc, ) accelerates the first rolling hash phase of hash-based CDC algorithms. SS-CDC uses AVX-512 registers to create multiple rolling-heads, i.e., calculating the rolling hash on bytes from multiple regions of the file independently and in parallel. Each rolling head maintains its own hash value and independently calculates the contributions of incoming and outgoing bytes.
To use vector instructions, they first collect the outgoing bytes for each head into a vector register $V_{1}$. Similarly, they collect all the incoming bytes into another vector register $V_{2}$. The hash values for each head are stored in a separate register $V_{3}$. SS-CDC removes the contributions of all outgoing bytes from the hash values with a single vector operation and adds the contributions of all incoming bytes with another. Whenever any of the hash values match the boundary condition (such as the lower k bits being equal to zero in Rabin-Karp chunking (lbfs, )), they mark the current position as a boundary candidate in a separate bitmap. In the second phase, they scan the bitmap using vector instructions to determine the actual boundaries among all candidates, taking into account the minimum and maximum chunk sizes.
@@
Second, to load incoming and outgoing bytes from different regions in the file, SS-CDC (sscdc, ) uses AVX gather instructions. To populate the candidate bitmap when boundary candidates are discovered, they use scatter instructions. These scatter and gather instructions are slow (scatter_gather_perf, ), limiting performance gains. Finally, scatter instructions are only available on processors supporting certain instruction sets (intelIntrinsicsGuide, ), limiting SS-CDC’s usage to a handful of the newest Intel and AMD processors (§2.3).
Figure 5. SS-CDC (sscdc, ) throughputs on randomized data with AVX-512 instructions
Figure 5 shows the chunking throughput obtained by running SS-CDC accelerated versions of CRC (SS-CRC) and Gear-based chunking (SS-Gear) (sscdc, ) against their native unaccelerated counterparts. This experiment used randomized data, an Intel Emerald Rapids machine described in §6 and AVX-512 instructions. We ran each algorithm with chunk sizes of 4164-16 KB. SS-CRC achieves 1.2 GB/s, a speedup of $2.58\times$ over CRC. Similarly, SS-Gear achieves 3.3 GB/s, a speedup of $3.13\times$ over Gear. These small speedups result from the challenges associated with hash-based algorithms that are described above.
@@ 4. VectorCDC Design
Figure 6. Accelerating Extreme Byte Search
Hashless CDC algorithms such as AE (ae, ), RAM (ram, ), and MAXP (maxp, ) slide windows over the source data to determine chunk boundaries. We identify two common processing patterns across all hashless CDC algorithms: the Extreme Byte Search and Range Scan. We accelerate each of these patterns using different vector-based techniques, which are discussed in detail below.
While we use the AVX-512 instruction set as an example to describe our acceleration techniques in this section, they can be implemented on any CPU with a vector instruction set supporting VMAX, VCMP and VMASK operations. SSE-128 and AVX-256 instruction sets (intelIntrinsicsGuide, ) fall under this umbrella, as do ARM and IBM processors with NEON-128 (armNeonOverview, ) and AltiVec / VSX-128 (ibm_power8, ) instructions, respectively. Thus, VectorCDC is compatible with a wide range of processors, unlike SS-CDC (sscdc, ), which relies on scatter instructions only available in AVX-512 instruction sets. Finally, while other minima/maxima-based hashless algorithms can also be accelerated using VectorCDC, their native versions are slower (ae, ; ram, ; maxp, ) than AE, RAM, and MAXP and have been omitted from the rest of our paper.
@@ 4.1. Tree-based Extreme Byte Search
Hashless CDC algorithms such as AE (ae, ), RAM (ram, ), and MAXP (maxp, ) all consist of a subsequence that identifies the extreme byte (maximum/minimum) in a fixed-size window. The size of this window depends upon the expected average chunk size and can be as large as 484-8KB. As this subsequence may need to be performed more than once per chunk, we propose accelerating it using a novel tree-based search approach. Let us consider the search for a maximum value using AVX-512 instructions (Figure 6). Note that the same method can be used with other vector instruction sets as well as to find minimum values.
We first divide the fixed-size window into smaller sub-regions, loading all the bytes into AVX-compatible m512i variables in Step 1. We load these bytes in a packed fashion i.e. each m512i variable contains 64 adjacent bytes. We then use vector mm512_max instructions to find the pairwise maximum among packed byte pairs (Step 2). For instance, among bytes ”E1” and ”21”, byte value ”E1” is the maximum. The resulting pairwise maximums are packed into a destination variable ($V_{5}$ in the figure).
@@ 4.2. Packed Scanning for Range Scans
Figure 7. Accelerating Range Scan
Hashless CDC algorithms also consist of a range scan subsequence, where bytes are serially compared against a target value. We propose to accelerate this scanning process using vector instructions. Let us consider a case where we compare bytes sequentially to see if they are greater than or equal to a target value (such as in RAM (ram, )). Figure 7 shows our proposal to accelerate this using packed scanning with AVX-512 instructions. Note that the same methods are applicable for other vector instruction sets as well.
We first load the target value (”F4” in Figure 7) into an AVX-compatible m512i variable $V_{1}$. We then pack 64 adjacent bytes from the scan region into another m512i variable $V_{2}$. We compare these 2 registers using mm512_cmpge vector compare instructions, which generate a 64-bit integer mask containing the comparison results. If this mask has a value greater than 0, a chunk boundary exists within the scanned 64 bytes. Its exact position is determined using the mask value. If the mask equals 0, no boundary exists within the scanned region and we proceed with loading the next 64 bytes into $V_{2}$ to repeat the process.
@@ 4.3. Putting it together: AE-Max, AE-Min, MAXP, and RAM
RAM (ram, ) first scans a fixed-size window at the beginning of the chunk to find a maximum value (Figure 2(c)). After this, it inserts a chunk boundary at the first byte outside the window, which is at least as large as the maximum valued byte (§2.1). With VectorCDC, we accelerate RAM as a combination of an Extreme Byte Search to find a maximum value, followed by a Range Scan with the GEQ comparator that compares this maximum value against bytes until a chunk boundary is found.
AE-Max (ae, ) scans for a byte larger than all the bytes before it i.e., a target byte (Figure 2(a)). Once found, a fixed-size window after this byte is scanned to determine the maximum valued byte within. If the target byte is larger than the maximum valued byte, a chunk boundary is inserted; otherwise, scanning continues for a new target byte (§2.1). With VectorCDC, we accelerate AE-Max as a combination of multiple Range Scans with the GT comparator to find target bytes, each followed by a single Extreme Byte Search for a maximum value.
AE-Min (ae, ) scans for a byte with lesser value than all those before it (§2.1). Once found, a fixed-size window after this byte is scanned to determine the minimum value within. If the target byte has a lesser value than the minimum valued byte, a chunk boundary is inserted; otherwise, scanning continues for a new target byte. Similar to AE-Max, we accelerate AE-Min as a combination of multiple Range Scans with the LT comparator to find target bytes, each followed by a single Extreme Byte Search for a minimum value.
Finally, MAXP (maxp, ) scans for a target local maxima that is exactly centered between two fixed-size windows (Figure 2(b)). A chunk boundary is inserted right after such a byte is found (§2.1). Thus, each chunk in MAXP can be represented as a combination of multiple Range Scans with the GT comparator, each followed by two Extreme Byte Searches for maximum values.

### v2

@@ 2.3. Vector Instructions
Vector instructions (Smith et al., 2000) are special Single-Instruction Multiple-Data (SIMD) instructions supported by most modern processors. They rely on special vector registers for their operations. These registers come in multiple sizes; depending on the width of the vector register they use, vector instructions can be classified into different families (Smith et al., 2000). The most common vector register sizes are 128 bits, 256 bits, and 512 bits, i.e., 16, 32, and 64 bytes wide.
@@
Vector instructions support various arithmetic operations (Intel, 2024), including pairwise addition, subtraction, multiplication, and maximum/minimum on packed values. Additionally, they support logical operations such as bitwise AND (&) and bitwise OR (|). They have been previously used to accelerate matrix multiplication (Hassan et al., 2018), sorting (Gueron and Krasnov, 2016), multimedia applications (Bocchino Jr and Adve, 2006), fluid simulations (Francés et al., 2014), hash tables (Böther et al., 2023), and relational databases (Dreseler et al., 2018). The supported vector instruction types and their relative performance vary across CPU manufacturers.
@@ 2.3.1. Intel and AMD
Vector instructions on x86 platforms can be classified into three families: SSE-128, AVX-256, and AVX-512 (Smith et al., 2000). SSE-128 instructions use 128-bit registers and have been supported by Intel and AMD processors since 1999 (Intel, ) and 2003 (Devices, 2003), respectively. AVX-256 instructions were introduced by Intel and AMD in 2011 (Intel, ), and use 256-bit registers. Finally, only a handful of the newest Intel and AMD processors, since 2017 (WikiChip, 2017) and 2022 (WikiChip, 2022), which have 512-bit wide vector registers support AVX-512 instructions.
@@ 2.3.2. ARM
ARM processors have supported NEON-128 instructions, an equivalent to SSE-128, since 2011 (ARM, 2013). Modern ARM processors also support vector widths of 256 bits and higher with the SVE/SVE2 instruction sets, which have been available since 2021 (Shi et al., 2025). These two instruction sets differ in the kinds of instructions supported. For instance, NEON-128 does not support native VMASK operations, which are used to create integer masks by extracting one out of every k bits in a vector register, while SVE / SVE2 does. This can lead to performance differences in applications that need the VMASK operation (Böther et al., 2023).
@@ 2.3.3. IBM Power
IBM’s Power (Sinharoy et al., 2015) architecture supports AltiVec / VSX-128 vector instructions (Miller et al., 2010), an equivalent to SSE-128, since the 1990s. This instruction set supports equivalents for most SSE instructions but lacks VMASK support.
@@ 3.1. Performance bottlenecks in data deduplication
Figure 3. Time taken for Data Chunking vs Fingerprinting while deduplicating randomized data with FastCDC (Xia et al., 2016b) and an 8 KB average chunk size.
Although both data chunking and fingerprinting have traditionally been considered the main bottlenecks in deduplication (Meyer and Bolosky, 2012), this has changed with the advent of new hashing algorithms and acceleration methods for fingerprinting. Figure 3 is a stacked bar plot showing the time taken by the data chunking and fingerprinting phases during deduplication. For fingerprinting, we use five different collision-resistant hashing algorithms (Rivest, 1992; Eastlake 3rd, 2001; Eastlake 3rd and Hansen, 2006; Appleby, 2011; Chi and Zhu, 2017). For data chunking, we use FastCDC (Xia et al., 2016b), the fastest unaccelerated chunking algorithm (§6.2), with an 8 KB average chunk size. We use 30 GB of randomized data and an Intel Emerald Rapids machine described in §6 for this experiment.
Figure 3 shows that fingerprinting and data chunking take nearly equal time with traditional collision-resistant hashing algorithms, such as SHA1 (Eastlake 3rd, 2001) and SHA-256 (Eastlake 3rd and Hansen, 2006). Fingerprinting takes longer than data chunking with other algorithms, such as MD5 (Rivest, 1992) and SHA-512 (Eastlake 3rd and Hansen, 2006). This indicates that both phases used to be the performance bottlenecks in deduplication, aligning with previous literature (Meyer and Bolosky, 2012).
However, recent research has introduced faster hashing algorithms such as MurmurHash3 (Appleby, 2011) and xxHash3 (xxHash, 2020) that generate 128-bit digests equivalent to MD5 (Appleby, 2016). These hashing algorithms are being used for fingerprinting (Tan et al., 2023) or as weak fingerprints followed by a byte-by-byte comparison (Ni et al., 2018). Figure 3 shows that fingerprinting takes significantly lower time than chunking with these new hashing algorithms, as they are $10\times-15\times$ faster than their counterparts on CPUs. Using GPUs can further accelerate fingerprinting speeds by up to $53\times$ (Suttisirikul and Uthayopas, 2012; Al-Kiswany et al., 2008).
Thus, as a result of its computationally intensive nature and position on the critical path, data chunking is a prime target for acceleration.
@@ 3.2. Accelerating hash-based algorithms with vector instructions
Figure 4. SS-CDC (Ni et al., 2019): Accelerating the rolling hash phase
To address the data chunking bottleneck, SS-CDC (Ni et al., 2019) proposed using AVX-512 instructions to accelerate hash-based CDC algorithms. They decouple the rolling hash and boundary detection phases, running the rolling hash on the entire source data to identify boundary candidates in the first phase, and determining boundaries sequentially in the second. This allows both stages to be independently accelerated with AVX instructions.
Figure 4 shows how SS-CDC (Ni et al., 2019) accelerates the first rolling hash phase of hash-based CDC algorithms. SS-CDC uses AVX-512 registers to create multiple rolling-heads (Head 1 - Head k), i.e., calculating the rolling hash on bytes from multiple regions of the file independently and in parallel. Each rolling head maintains its own hash value in the hash value register and independently calculates the contributions of incoming and outgoing bytes.
To use vector instructions, they first collect the outgoing bytes for each head into a vector register $V_{1}$. Similarly, they collect all the incoming bytes into another vector register $V_{2}$. This is shown in Step 1 in Figure 4 and uses gather instructions. The hash values for each head are stored in a separate register $V_{3}$. In Step 2, SS-CDC removes the contributions of all outgoing bytes from the hash values with a single vector operation and adds the contributions of all incoming bytes with another.
Step 3 compares all the hash values against the pre-specified boundary condition (such as the lower x bits being equal to zero in Rabin-Karp chunking (Muthitacharoen et al., 2001)). Whenever any of the hash values match the boundary condition, they mark the current position as a boundary candidate in a separate bitmap in Step 4 using scatter instructions. This rolling hash phase is run on the entire incoming data stream/file. In the second phase, they scan the bitmap using vector instructions to determine the actual boundaries among all candidates, taking into account the minimum and maximum chunk sizes.
@@
Second, to load incoming and outgoing bytes from different regions in the file, SS-CDC (Ni et al., 2019) uses AVX gather instructions. To populate the candidate bitmap when boundary candidates are discovered, they use scatter instructions. These scatter and gather instructions are slow (Lavin et al., 2021), limiting performance gains. Finally, scatter instructions are only available on processors supporting certain instruction sets (Intel, 2024), limiting SS-CDC’s usage to a handful of the newest Intel and AMD processors (§2.3).
Figure 5. SS-CDC (Ni et al., 2019) throughputs on randomized data with AVX-512 instructions
Figure 5 shows the chunking throughput obtained by running SS-CDC accelerated versions of CRC (SS-CRC) and Gear-based chunking (SS-Gear) (Ni et al., 2019) against their native unaccelerated counterparts. This experiment used randomized data, an Intel Emerald Rapids machine described in §6 and AVX-512 instructions. We ran each algorithm with chunk sizes of 4164-16 KB. SS-CRC achieves 1.2 GB/s, a speedup of $2.58\times$ over CRC. Similarly, SS-Gear achieves 3.3 GB/s, a speedup of $3.13\times$ over Gear. These small speedups result from the challenges associated with hash-based algorithms that are described above.
@@ 4. VectorCDC Design
Hashless CDC algorithms such as AE (Zhang et al., 2015), RAM (Widodo et al., 2017), and MAXP (Bjørner et al., 2010) slide windows over the source data to determine chunk boundaries. We identify two common processing patterns across all hashless CDC algorithms: the Extreme Byte Search and Range Scan. We accelerate each of these patterns using different vector-based techniques, which are discussed in detail below.
While we use the AVX-512 instruction set as an example to describe our acceleration techniques in this section, they can be implemented on any CPU with a vector instruction set supporting VMAX, VCMP and VMASK operations. SSE-128 and AVX-256 instruction sets (Intel, 2024) fall under this umbrella, as do ARM and IBM processors with NEON-128 (ARM, 2013) and AltiVec / VSX-128 (Sinharoy et al., 2015) instructions, respectively. Thus, VectorCDC is compatible with a wide range of processors, unlike SS-CDC (Ni et al., 2019), which relies on scatter instructions only available in AVX-512 instruction sets. Finally, while other minima/maxima-based hashless algorithms can also be accelerated using VectorCDC, their native versions are slower (Zhang et al., 2015; Widodo et al., 2017; Bjørner et al., 2010) than AE, RAM, and MAXP and have been omitted from the rest of our paper.
Figure 6. Accelerating Extreme Byte Searches. Note that the byte values shown are in hexadecimal format.
@@ 4.1. Tree-based Extreme Byte Search
Hashless CDC algorithms such as AE (Zhang et al., 2015), RAM (Widodo et al., 2017), and MAXP (Bjørner et al., 2010) all consist of a subsequence that identifies the extreme byte (maximum/minimum) in a fixed-size window. The size of this window depends upon the expected average chunk size and can be as large as 484-8KB. As this subsequence may need to be performed more than once per chunk, we propose accelerating it using a novel tree-based search approach. Let us consider the search for a maximum value using AVX-512 instructions (Figure 6). Note that the same method can be used with other vector instruction sets as well as to find minimum values.
We first divide the fixed-size window into smaller sub-regions, loading all the bytes into AVX-compatible m512i variables in Step 1. We load these bytes in a packed fashion i.e. each m512i variable contains 64 adjacent bytes. We then use vector mm512_max instructions to find the pairwise maximum among packed byte pairs (Step 2). For instance, among the bytes 0xE1 and 0x21, byte value 0xE1 is the maximum. The resulting pairwise maximums are packed into a destination variable ($V_{5}$ in the figure).
@@ 4.2. Packed Scanning for Range Scans
Figure 7. Accelerating Range Scans. Note that the byte values shown are in hexadecimal format.
Hashless CDC algorithms also consist of a range scan subsequence, where bytes are serially compared against a target value. We propose to accelerate this scanning process using vector instructions. Let us consider a case where we compare bytes sequentially to see if they are greater than or equal to a target value (such as in RAM (Widodo et al., 2017)). Figure 7 shows our proposal to accelerate this using packed scanning with AVX-512 instructions. Note that the same methods are applicable for other vector instruction sets as well.
We first load the target value (0xF4 in Figure 7) into an AVX-compatible m512i variable $V_{1}$. We then pack 64 adjacent bytes from the scan region into another m512i variable $V_{2}$. We compare these 2 registers using mm512_cmpge vector compare instructions, which generate a 64-bit integer mask containing the comparison results. If this mask has a value greater than 0, a chunk boundary exists within the scanned 64 bytes. Its exact position is determined using the mask value. If the mask equals 0, no boundary exists within the scanned region and we proceed with loading the next 64 bytes into $V_{2}$ to repeat the process.
@@ 4.3. Putting it together: AE-Max, AE-Min, MAXP, and RAM
RAM (Widodo et al., 2017) first scans a fixed-size window at the beginning of the chunk to find a maximum value (Figure 2(c)). After this, it inserts a chunk boundary at the first byte outside the window, which is at least as large as the maximum valued byte (§2.1). With VectorCDC, we accelerate RAM as a combination of an Extreme Byte Search to find a maximum value, followed by a Range Scan with the GEQ comparator that compares this maximum value against bytes until a chunk boundary is found.
AE-Max (Zhang et al., 2015) scans for a byte larger than all the bytes before it i.e., a target byte (Figure 2(a)). Once found, a fixed-size window after this byte is scanned to determine the maximum valued byte within. If the target byte is larger than the maximum valued byte, a chunk boundary is inserted; otherwise, scanning continues for a new target byte (§2.1). With VectorCDC, we accelerate AE-Max as a combination of multiple Range Scans with the GT comparator to find target bytes, each followed by a single Extreme Byte Search for a maximum value.
AE-Min (Zhang et al., 2015) scans for a byte with lesser value than all those before it (§2.1). Once found, a fixed-size window after this byte is scanned to determine the minimum value within. If the target byte has a lesser value than the minimum valued byte, a chunk boundary is inserted; otherwise, scanning continues for a new target byte. Similar to AE-Max, we accelerate AE-Min as a combination of multiple Range Scans with the LT comparator to find target bytes, each followed by a single Extreme Byte Search for a minimum value.
Finally, MAXP (Bjørner et al., 2010) scans for a target local maxima that is exactly centered between two fixed-size windows (Figure 2(b)). A chunk boundary is inserted right after such a byte is found (§2.1). Thus, each chunk in MAXP can be represented as a combination of multiple Range Scans with the GT comparator, each followed by two Extreme Byte Searches for maximum values.